OpenStudy (anonymous):

2-x/4 <= -3 how can I solve it

OpenStudy (anonymous):

send x/4 to right and bring -3 to left 2+3<=x/4 x/4>=5 z>=5*4

OpenStudy (ciarán95):

We take the -x/4 term to the right-hand side of the inequality, and take the -3 on the right-hand side over to the left-hand side. In doing so, we change the signs on both. So, 2 - (x/4) <= -3 becomes 2 + 3 <= x/4 (we're simply treating the <= as if it were just an = sign). This gives us 5 <= x/4. To get x on its own we need to multiply each side of the inequality by 4. As 4 is a positive number we do not need to reverse the inequality sign. We would only need to this if it was a negative value. So, we get (4)(5) <= x, which we can rewrite as x >= 20.

OpenStudy (anonymous):

\(\Large 2-\frac{x}{4} \le -3 \Leftrightarrow \frac{x}{4} \ge 2+3 \Rightarrow x\ge 4\times5 \Rightarrow x\ge20\)